If f(x)=(x+1)/x, x ≠ 0 and g(x)= 2x+3, find g°f(x)

1
Expert's answer
2022-05-03T09:29:01-0400
"g\circ f(x)=2(\dfrac{x+1}{x})+3=\dfrac{5x+2}{x},x\not=0"